OpenStudy (anonymous):

lim(x->0) (sin^2(2 x))\/x^2

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ \sin ^{2}(2x) }{x^2 }\]

terenzreignz (terenzreignz):

Seems too much to hope, but is L'Hôpital allowed?

OpenStudy (anonymous):

Sadly no :(

OpenStudy (anonymous):

Can use Sinx/x = 1 and 1-cosx/x = 0

terenzreignz (terenzreignz):

Well in that case, why don't we try and decompose the function into something that resembles those two forms?

terenzreignz (terenzreignz):

How about this: \[\Large \frac{\sin^2(2x)}{x^2}= \frac{\sin(2x)}{x}\cdot\frac{\sin(2x)}{x}\]

OpenStudy (anonymous):

I was thinking \[\frac{ (Sin2x)(Sin2x) }{x^2 }*\frac{ 2 }{ 2 }\] Does this make sense?

OpenStudy (anonymous):

Then I come up with 2 as Sin2x/2x =1

terenzreignz (terenzreignz):

But you need two of them.

OpenStudy (anonymous):

Ohhhh

terenzreignz (terenzreignz):

\[\Large \frac{\sin^2(2x)}{x^2}= \frac{\sin(2x)}{\color{red}x}\cdot\frac{\sin(2x)}{\color{green}x}\]

OpenStudy (anonymous):

\[\frac{ \sin^2(2x) }{ x^2 }*\frac{ 2 }{ 2 }*\frac{ 2 }{ 2 } = 4\]

terenzreignz (terenzreignz):

Bingo :P

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

Nice <3

terenzreignz (terenzreignz):

Though, allow me this one little nitpick: \[\Large \color{red}{\lim_{x\rightarrow0}}\frac{ \sin^2(2x) }{ x^2 }*\frac{ 2 }{ 2 }*\frac{ 2 }{ 2 } = 4\]

OpenStudy (anonymous):

Yup, perfect.